Electric field and electric potential at a point 3cm away from a point charge is E and V respectively. What is the electric field and electric potential at 9 cm away from the same charge.


1
Expert's answer
2021-05-05T10:54:06-0400
"E'=\frac{3^2}{9^2}E=\frac{1}{9}E"

"V'=\frac{3}{9}V=\frac{1}{3}V"